OpenStudy (anonymous):

q. in the hydrogen atom the energy of an electron in first bojrs orbit is 13.12*10raised to the power 5 joule mol-1?what is the energy required for its excitation to bohrs second orbit?

OpenStudy (mos1635):

there is a formula for Energyof orbit n (n=1,2,3,...)

OpenStudy (mos1635):



OpenStudy (mos1635):



OpenStudy (anonymous):

please provide me solution with final answer plzz as i dont know anyhng about this

OpenStudy (mos1635):

the formula for radious is \[r _{n}= n ^{2}\times r _{1}\]

OpenStudy (mos1635):

do you know the formula for energy?? (look it up)

OpenStudy (anonymous):

then what to do next

OpenStudy (anonymous):

then what to do next

OpenStudy (anonymous):

ya i know the formula of energyy then me aftr that wht to doo next after that

OpenStudy (anonymous):

ya i know the formula of energyy then me aftr that wht to doo next after that

OpenStudy (mos1635):

find E2

OpenStudy (anonymous):

please tell me whole solution

OpenStudy (mos1635):

\[E _{2}=\frac{ E _{1} }{ n ^{2} }\]

OpenStudy (mos1635):

ΔE=E1-E2

OpenStudy (mos1635):

INITIAL GIVEN ENERGY E1 Ι ΤΗΙΝΚ ΙT should be negative E1=-13.12*10^5 if so ΔΕ=E2-E1

OpenStudy (anonymous):

thennnnn

OpenStudy (anonymous):

wht is next after this

OpenStudy (anonymous):

plzzzzzzzzz telll

OpenStudy (abb0t):

Giving direct answers would violate CoC. And most importantly you wouldn't learn anything. http://openstudy.com/code-of-conduct

OpenStudy (preetha):

Kapuria - we need to see your effort as well.

OpenStudy (anonymous):

i think answer is 9.84*10raised to the power 5 joule

OpenStudy (mos1635):

that is correct. (please diascase with your teacher my coment regarding the (-) to initial given energy.....

OpenStudy (mos1635):

joule/mole

OpenStudy (anonymous):

kkkkkkk listen apne jo muujhe vatya unki values bhi batdo E1=-13.12*10^5 if so ΔΕ=E2-E1 then wht is the next step plzzzzzzzzz

OpenStudy (anonymous):

PLZZZ REQUEST